OpenStudy (anonymous):

Write the equation of a line containing the point (2,3) and that has a slope 0f -3? hope to get answer and explain

OpenStudy (anonymous):

y=-3x+3

OpenStudy (anonymous):

What is f?

OpenStudy (anonymous):

I think she means that the equation has a slope of -3

OpenStudy (anonymous):

yes typing mistake

OpenStudy (anonymous):

Let me check that real quick actually

OpenStudy (anonymous):

ok thats actually wrong After graphing it I saw that the y intercept is actually 9 3=-3(2)+b 3=-6+b 9=b Slope-intercept form y=-3x+9

OpenStudy (anonymous):

thank i got it

OpenStudy (anonymous):

lol Ya sorry about that